# Completions of Kleene’s second model

### v3

@@ Abstract.
We investigate completions of partial combinatory algebras (pcas), in particular of Kleene’s second model $\mathcal{K}_{2}$ and generalizations thereof. We consider weak and strong notions of embeddability and completion that have been studied before. By a result of Klop it is known that not every pca has a strong completion. The study of completions of $\mathcal{K}_{2}$ has as corollaries that weak and strong embeddings are different, and that every countable pca has a weak completion. We then consider generalizations of $\mathcal{K}_{2}$ for larger cardinals, and use these to show that it is consistent that every pca has a weak completion.
@@
partial combinatory algebra, completions
@@ 1. Introduction
A completion of a pca is an embedding into a (total) combinatory algebra. Here one can study weak and strong kinds of completion, see Definition 2.1. In the strong version, the combinators $k$ and $s$ that every pca possesses are considered as part of the signature, and in the weak version they are not. Both kinds of embedding have been studied in the literature. The question (posed by Barendregt, Mitschke, and Scott) whether every pca has a strong completion was answered in the negative by Klop [11], see also Bethke, Klop, and de Vrijer [5]. On the other hand, $\mathcal{K}_{2}$ does have strong completions, see Theorem 3.3 below. Weak completions and embeddings were studied for example in Bethke [3], Asperti and Ciabattoni [1], Shafer and Terwijn [19], Zoethout [22], Golov and Terwijn [8], and more recently in Fokina and Terwijn [7]. There is an even weaker notion of embeddability of pcas introduced by Longley, see [13] and [14], that is useful in the study of realizability. We will not consider this notion here, but simply note that the results about weak completions obtained here are the strongest possible.
The study of completions of $\mathcal{K}_{2}$ has as corollaries that weak and strong embeddings are different (Theorem 4.1), and that every countable pca has a weak completion (Corollary 2.2), which contrasts the result of Klop for strong completions. We then consider generalizations of $\mathcal{K}_{2}$ for larger cardinals, and use these to show that it is consistent that every pca has a weak completion (Theorem 6.2).
@@ 2. Preliminaries
There exists pcas that do not have a strong completion, see Klop [11] and Bethke, Klop, and de Vrijer [5]. On the other hand, we can see that every countable pca has a weak completion as follows. In Golov and Terwijn [8, Corollary 6.2] it was shown that every countable pca is weakly embeddable into $\mathcal{K}_{2}$. Now $\mathcal{K}_{2}$ is weakly completable, as it weakly embeds, for example, into Scott’s graph model $\mathcal{G}$ (see Scott [18] for the definition, and [8] for the embedding), which is a total combinatory algebra related to enumeration reducibility. Hence we obtain the following corollary:
Corollary 2.2.
All countable pcas have a weak completion.
In Theorem 6.2 below we prove that it is consistent that all pcas have a weak completion. We first show, in the next section, that $\mathcal{K}_{2}$ and $\mathcal{K}_{2}^{\mathrm{eff}}$ have strong completions. Note that none of these completions is computable, as (total) combinatory algebras are never computable. The complexity of completions was studied in [21].
@@ 3. Unique head normal forms
4. Weak versus strong embeddings
The strong completability of $\mathcal{K}_{2}$ has as consequence that weak and strong embeddings are different.
Theorem 4.1.
Weak and strong embeddability and completions are different: There exists a pca that is weakly completable, but not strongly completable.
Proof.
Let $\mathcal{A}$ be the example from Bethke et al. [5] of a pca that is not strongly completable. We note that $\mathcal{A}$ is a countable pca. By Golov and Terwijn [8], there exists a weak embedding $\mathcal{A}\hookrightarrow\mathcal{K}_{2}$. Since by Theorem 3.3, $\mathcal{K}_{2}$ is strongly completable, it follows immediately that the embedding $\mathcal{A}\hookrightarrow\mathcal{K}_{2}$ cannot be a strong one, so that weak and strong embeddability must be different. Also, the strong completion of $\mathcal{K}_{2}$ is automatically a weak completion of $\mathcal{A}$, so that $\mathcal{A}$ has a weak, but not a strong completion. ∎
@@ 5. $\mathcal{K}_{2}$ generalized
We will make use of the following generalization of $\mathcal{K}_{2}$, taken from van Oosten [17, 1.4.4].
@@
Following Kunen [12, p34], we define $\kappa^{<\kappa}=\bigcup\{\kappa^{\lambda}:\lambda<\kappa\}$. For a function $\sigma\in\kappa^{<\kappa}$, we let $\mathrm{dom}(\sigma)$ denote the largest $\lambda<\kappa$ on which it is defined. By assuming $\kappa^{<\kappa}=\kappa$, we have a bijective coding function $\langle\cdot\rangle:\kappa^{<\kappa}\to\kappa$. For $n\in\kappa$ and $\sigma$ an element of either $\kappa^{<\kappa}$ or $\kappa^{\kappa}$, we let $n\mbox{}^{\smallfrown}\sigma$ denote the function defined by
@@
Now if we can also make $F$ invertible, it will follow that $s$ such that $s\alpha\beta=F(\alpha,\beta)$, is also invertible, and hence satisfies Barendregt’s axiom (5). However, the definition of $F$ obtained from unwinding (8) does not automatically give that $F$ is invertible, so we need to do a bit of extra work for this. We show that, starting with any $F$ satisfying (8), we can convert $F$ to a function that is in addition invertible. The idea is that if $(F(\alpha,\beta)\gamma)(n)$ is defined for any $\gamma\in\kappa^{\kappa}$ and $n\in\kappa$, then the coding of Definition 5.1 leaves ample room to code $\alpha$ and $\beta$ in the values of $F(\alpha,\beta)$ that do not matter. The only trouble is when there are no defined values at all, in which case we have to resort to another solution.
@@
Case 2. $(F(\alpha,\beta)\gamma)(n)$ is undefined for all $\gamma\in\kappa^{\kappa}$ and $n\in\kappa$. In this case we have by Definition 5.1 that $F(\alpha,\beta)(n)$ is equal to 00 or a limit for all $n\in\kappa$. The effect of $F(\alpha,\beta)$ does not change if we change the value from one limit or 00 to another. W.l.o.g. $\kappa>\omega$, since for $\kappa=\omega$ we already have a proof (Lemma 3.2). If $\kappa>\omega$ we have the two values $0,\omega<\kappa$ that we can use for coding $\alpha$ and $\beta$. Again we can code these as one sequence $\alpha\oplus\beta\in\kappa^{\kappa}$ Using a bijection $\kappa^{\kappa}\leftrightarrow 2^{\kappa}$ we may assume $\alpha\oplus\beta\in 2^{\kappa}$, and code it using the two values 00 and $\omega$ by
@@
This concludes the proof that we can make $F$ invertible without changing its behavior. As we already noted, the resulting combinator $s$ then satisfies Barendregt’s axiom (5). In particular, all the hnfs $s\alpha\beta$ are unique. Since $\alpha$ can be retrieved from $s\alpha\beta$ for any $\beta$, it follows that $s\alpha$ is also invertible, and hence the hnfs $s\alpha$ are also unique. It follows from the properties of $k$ that these are also different from all $k\alpha$, so that indeed all hnfs are different. ∎
@@
For $\mathcal{K}_{2}$ (i.e. for $\kappa=\omega$), we gave two arguments for its strong completability in Section 3. Apart from the argument using Lemma 3.2 about unique hnfs we had an easier option using van Oosten’s model $\mathcal{B}$. It would be interesting to generalize this to $\mathcal{B}^{\kappa}$ for all $\kappa$ with $\kappa^{<\kappa}=\kappa$ as above. It is not immediately clear how to define $\mathcal{B}^{\kappa}$. Simply restricting $\mathcal{K}_{2}^{\kappa}$ to partial functions does not work, as this makes use of the machine model for $\mathcal{K}_{2}$ that is not available for $\mathcal{K}_{2}^{\kappa}$. Also, the coding from Definition 5.1 is not suitable for $\mathcal{B}$, as is discussed in Longley and Normann [14, p77], because this coding does not allow to jump over undefined values.
The definition of $\mathcal{K}_{2}^{\kappa}$ uses the property $\kappa^{<\kappa}=\kappa$, needed for the coding $\langle\cdot\rangle:\kappa^{<\kappa}\to\kappa$. It is known that inaccessible cardinals have this property, cf. Jech [9, p50], but the existence of inaccessibles cannot be proved in $\mathrm{ZFC}$ (cf. Kunen [12]). For $\kappa=2^{\omega}$, the property $\kappa^{<\kappa}=\kappa$ is independent. [^2] One can easily check that the statement is true under CH and false e.g. when $2^{\omega}=\omega_{2}$ and $2^{\omega_{1}}=\omega_{3}$. In Theorem 6.2 we will use $\mathcal{K}_{2}^{\kappa}$ to show that it is consistent that all pcas have a weak completion. To guarantee existence of $\mathcal{K}_{2}^{\kappa}$ for arbitrary large $\kappa$ we can for example use the generalized continuum hypothesis $\mathrm{GCH}$, which says that $2^{\omega_{\alpha}}=\omega_{\alpha+1}$ for every ordinal $\alpha$. (This is consistent by Gödels famous result that $\mathrm{GCH}$ holds in $\mathrm{L}$.) Assuming $\mathrm{GCH}$ we have
@@
6. Weak completability of all pcas
As mentioned above, Klop [11, 5] showed that not every pca has a strong completion. Asperti and Ciabattoni [1] proved that if Barendregt’s axiom (5) holds in a pca then it has a weak completion. See also the discussion about this in Bethke et al. [5, p501].
Note that Barendregt’s axiom does not hold for the example from Bethke et al. [5] of a pca without strong completions. Indeed, on p486 loc. cit. we have $a$ and $b$ distinct such that $s(sk)a=s(sk)b$, which violates Barendregt’s axiom. So we cannot conclude from the Asperti-Ciabattoni result that the example has a weak completion. But since the example constructed in [5] is countable, it follows from Corollary 2.2 that it does have a weak completion.
Having seen that every countable pca weakly embeds into $\mathcal{K}_{2}$, and hence has a weak completion, we want to extend this result to larger cardinalities. Now we do have the large pcas $\mathcal{K}_{2}^{\kappa}$, which is a pca on $\kappa^{\kappa}$, but as discussed in Section 5 the definition of this is contingent on the condition $\kappa^{<\kappa}=\kappa$, which implies that for $\kappa>\omega$ at best we know that their existence is consistent with $\mathrm{ZFC}$ (using $\mathrm{GCH}$). On the basis of this we can show that every pca has a weak completion.
@@ 6. Weak completability of all pcas
Theorem 6.2.
It is consistent with $\mathrm{ZFC}$ that all pcas are weakly completable.
Proof.
Assume $\mathrm{GCH}$. By Theorem 6.1 there exists $\kappa$ such that $\mathcal{A}$ weakly embeds into $\mathcal{K}_{2}^{\kappa}$. By Theorem 5.4 $\mathcal{K}_{2}^{\kappa}$ has a strong completion, which is then also a weak completion of $\mathcal{A}$. ∎
As a corollary to Theorem 6.2 we see that it is not possible to prove in $\mathrm{ZFC}$ an analog of Klop’s result (that there exist pcas without strong completions) for weak completions. An obvious question that remains at this point is: Is it possible to prove the statement of Theorem 6.2 absolutely?

### v5

@@ Abstract.
We investigate completions of partial combinatory algebras (pcas), in particular of Kleene’s second model $\mathcal{K}_{2}$ and generalizations thereof. We consider weak and strong notions of embeddability and completion that have been studied before in the literature. It is known that every countable pca can be weakly embedded into $\mathcal{K}_{2}$, and we generalize this to arbitrary cardinalities by considering generalizations of $\mathcal{K}_{2}$ for larger cardinals. This emphasizes the central role of $\mathcal{K}_{2}$ in the study of pcas. We also show that $\mathcal{K}_{2}$ and its generalizations have strong completions.
@@
partial combinatory algebra, embeddings, completions
@@ 1. Introduction
A completion of a pca is an embedding into a (total) combinatory algebra. Here one can study weak and strong kinds of completion, see Definition 2.1. In the strong version, the combinators $k$ and $s$ that every pca possesses are considered as part of the signature, and in the weak version they are not. Both kinds of embedding have been studied in the literature. The question (posed by Barendregt, Mitschke, and Scott) whether every pca has a strong completion was answered in the negative by Klop [13], see also Bethke, Klop, and de Vrijer [5]. On the other hand, $\mathcal{K}_{2}$ does have strong completions, see Theorem 3.3 below. Weak completions and embeddings were studied for example in Bethke [3], Asperti and Ciabattoni [1], Shafer and Terwijn [22], Zoethout [25], Golov and Terwijn [10], and more recently in Fokina and Terwijn [9]. It follows from the main result in Engeler [6] that every pca has a weak completion. Hence fixing the combinators $k$ and $s$ or not in embedding results makes a big difference. There is an even weaker notion of embeddability of pcas introduced by Longley, see [15] and [16], that is useful in the study of realizability. We will not consider this notion here, but simply note that the results about weak completions obtained here are the strongest possible. Extensions of Kleene’s second model containing specific functions were studied in van Oosten and Voorneveld [20].
It is known that every countable pca is weakly embeddable into $\mathcal{K}_{2}$. This is optimal, since this does not hold for strong embeddings. We generalize this to larger cardinalities by considering a generalization $\mathcal{K}_{2}^{\kappa}$ for larger cardinals $\kappa$ (Section 4). We show that $\mathcal{K}_{2}$ and $\mathcal{K}_{2}^{\kappa}$ have strong completions (Theorems 3.3 and 4.4), and prove an embedding result for $\mathcal{K}_{2}^{\kappa}$ (Theorem 5.1).
@@ 2. Preliminaries
There exists pcas that do not have a strong completion, see Klop [13] and Bethke, Klop, and de Vrijer [5]. On the other hand, it follows from Engeler [6] that every pca has a weak completion.111 The main result in [6] is stated only for structures with one binary operator, but on p390 it is mentioned that the representation theorem can be generalized to multiple and also to partial operations, so that in particular it also holds for pcas. This claim is further substantiated in the supplementary note [7]. So in particular we see that weak and strong embeddings and completions are different.
We can see that every countable pca has a weak completion as follows. In Golov and Terwijn [10, Corollary 6.2] it was shown that every countable pca is weakly embeddable into $\mathcal{K}_{2}$. Now $\mathcal{K}_{2}$ is weakly completable, as it weakly embeds, for example, into Scott’s graph model $\mathcal{G}$ (see Scott [21] for the definition, and [10] for the embedding), which is a total combinatory algebra related to enumeration reducibility. Hence every countable pcas has a weak completion. Note that this includes the counterexample for strong completions from Bethke et al. [5], which is countable.
In the next section we show that $\mathcal{K}_{2}$ and $\mathcal{K}_{2}^{\mathrm{eff}}$ have strong completions. Note that none of these completions is computable, as (total) combinatory algebras are never computable. The complexity of completions was studied in [24].
@@ 4. $\mathcal{K}_{2}$ generalized
We will make use of the following generalization of $\mathcal{K}_{2}$, taken from van Oosten [19, 1.4.4].
@@
Following Kunen [14, p34], we define $\kappa^{<\kappa}=\bigcup\{\kappa^{\lambda}:\lambda<\kappa\}$. For a function $\sigma\in\kappa^{<\kappa}$, we let $\mathrm{dom}(\sigma)$ denote the largest $\lambda<\kappa$ on which it is defined. By assuming $\kappa^{<\kappa}=\kappa$, we have a bijective coding function $\langle\cdot\rangle:\kappa^{<\kappa}\to\kappa$. For $n\in\kappa$ and $\sigma$ an element of either $\kappa^{<\kappa}$ or $\kappa^{\kappa}$, we let $n\mbox{}^{\smallfrown}\sigma$ denote the function defined by
@@
Now if we can also make $F$ injective, it will follow that $s$ such that $s\alpha\beta=F(\alpha,\beta)$, is also injective, and hence satisfies Barendregt’s axiom (5). However, the definition of $F$ obtained from unwinding (8) does not automatically give that $F$ is injective, so we need to do a bit of extra work for this. We show that, starting with any $F$ satisfying (8), we can convert $F$ to a function that is in addition injective. The idea is that if $(F(\alpha,\beta)\gamma)(n)$ is defined for just any $\gamma\in\kappa^{\kappa}$ and $n\in\kappa$, then the coding of Definition 4.1 leaves ample room to code $\alpha$ and $\beta$ in the values of $F(\alpha,\beta)$ that do not matter. The only trouble is when there are no defined values at all, in which case we have to resort to another solution.
@@
Case 2. $(F(\alpha,\beta)\gamma)(n)$ is undefined for all $\gamma\in\kappa^{\kappa}$ and $n\in\kappa$. In this case we have by Definition 4.1 that $F(\alpha,\beta)(n)$ is equal to 00 or a limit for all $n\in\kappa$. The effect of $F(\alpha,\beta)$ does not change if we change the value from one limit or 00 to another. W.l.o.g. $\kappa>\omega$, since for $\kappa=\omega$ we already have a proof (Lemma 3.2). If $\kappa>\omega$ we have the two values $0,\omega<\kappa$ that we can use for coding $\alpha$ and $\beta$. Again we can code these as one sequence $\alpha\oplus\beta\in\kappa^{\kappa}$. Using a bijection $\kappa^{\kappa}\leftrightarrow 2^{\kappa}$ we may assume $\alpha\oplus\beta\in 2^{\kappa}$, and code it using the two values 00 and $\omega$ by
@@
Now since the adapted version of $F$ still has to be continuous, we cannot make the case distinction above separately for every $\alpha$ and $\beta$: For some $\alpha$ and $\beta$ where case 1 applies, there may be close $\alpha^{\prime}$ and $\beta^{\prime}$ where case 2 applies, and since the coding for $\alpha^{\prime}$ and $\beta^{\prime}$ starts from the beginning, we have no choice but to do the same for $\alpha$ and $\beta$. So we define the adaptation of $F(\alpha,\beta)(\langle\sigma\rangle))$, with $\sigma\in\kappa^{<\kappa}$, by switching between the two cases dynamically. As long as we have not found a point yet where case 1 applies, we follow the coding of case 2, and otherwise we switch to the coding of case 1. Specifically, if for every $\sigma^{\prime}\sqsubseteq\sigma$, $F(\alpha,\beta)(\langle\sigma^{\prime}\rangle))$ is 00 or a limit, we apply the coding of case 2. Otherwise, if $\sigma^{\prime}\sqsubseteq\sigma$ is the minimal initial such that $F(\alpha,\beta)(\langle\sigma^{\prime}\rangle))$ is a successor, we apply from that point onwards the coding of case 1.
This concludes the proof that we can make $F$ injective without changing its behavior. As we already noted, the resulting combinator $s$ then satisfies Barendregt’s axiom (5). In particular, all the hnfs $s\alpha\beta$ are unique. Since $\alpha$ can be retrieved from $s\alpha\beta$ for any $\beta$, it follows that $s\alpha$ is also injective, and hence the hnfs $s\alpha$ are also unique. It follows from the properties of $k$ that these are also different from all $k\alpha$, so that indeed all hnfs are different. ∎
@@
For $\mathcal{K}_{2}$ (i.e. for $\kappa=\omega$), we gave two arguments for its strong completability in Section 3. Apart from the argument using Lemma 3.2 about unique hnfs we had an easier option using van Oosten’s model $\mathcal{B}$. It would be interesting to generalize this to $\mathcal{B}^{\kappa}$ for all $\kappa$ with $\kappa^{<\kappa}=\kappa$ as above. It is not immediately clear how to define $\mathcal{B}^{\kappa}$. Simply restricting $\mathcal{K}_{2}^{\kappa}$ to partial functions does not work, as this makes use of the machine model for $\mathcal{K}_{2}$ that is not available for $\mathcal{K}_{2}^{\kappa}$. Also, the coding from Definition 4.1 is not suitable for $\mathcal{B}$, as is discussed in Longley and Normann [16, p77], because this coding does not allow to jump over undefined values.
The definition of $\mathcal{K}_{2}^{\kappa}$ uses the property $\kappa^{<\kappa}=\kappa$, needed for the coding $\langle\cdot\rangle:\kappa^{<\kappa}\to\kappa$. It is known that inaccessible cardinals have this property, cf. Jech [11, p50], but the existence of inaccessibles cannot be proved in $\mathrm{ZFC}$ (cf. Kunen [14]). For $\kappa=2^{\omega}$, the property $\kappa^{<\kappa}=\kappa$ is independent. [^3] One can easily check that the statement is true under CH and false e.g. when $2^{\omega}=\omega_{2}$ and $2^{\omega_{1}}=\omega_{3}$. To guarantee existence of $\mathcal{K}_{2}^{\kappa}$ for arbitrary large $\kappa$ we can for example use the generalized continuum hypothesis $\mathrm{GCH}$, which says that $2^{\omega_{\alpha}}=\omega_{\alpha+1}$ for every ordinal $\alpha$. (This is consistent by Gödels famous result that $\mathrm{GCH}$ holds in $\mathrm{L}$.) Assuming $\mathrm{GCH}$ we have
@@
5. Embeddings into $\mathcal{K}_{2}^{\kappa}$
Having seen that every countable pca weakly embeds into $\mathcal{K}_{2}$, we want to extend this result to larger cardinalities. Now we do have the large pcas $\mathcal{K}_{2}^{\kappa}$, which is a pca on $\kappa^{\kappa}$, but as discussed in Section 4 the definition of this is contingent on the condition $\kappa^{<\kappa}=\kappa$, which implies that for $\kappa>\omega$ at best we know that their existence is consistent with $\mathrm{ZFC}$ (using $\mathrm{GCH}$).
@@ 5. Embeddings into $\mathcal{K}_{2}^{\kappa}$
6. Concluding remarks on completions
Asperti and Ciabattoni [1] proved that if Barendregt’s axiom (5) holds in a pca then it has a weak completion. (See also the discussion about this in Bethke et al. [5, p501].) By Engeler [6] (see footnote 1 above), the condition about Barendregt’s axiom is not needed. Note that Barendregt’s axiom does not hold for the example from Bethke et al. [5] of a pca without strong completions. Indeed, on p486 loc. cit. we have $a$ and $b$ distinct such that $s(sk)a=s(sk)b$, which violates Barendregt’s axiom. But since the example constructed in [5] is countable, and every countable pca weakly embeds into $\mathcal{K}_{2}$, it also follows from the strong completability of $\mathcal{K}_{2}$ that it has a weak completion.
Theorem 5.1 shows that, under GCH, every pca weakly embeds into a $\mathcal{K}_{2}^{\kappa}$ for suitably large $\kappa$. Since each $\mathcal{K}_{2}^{\kappa}$ is strongly completable, this has as a corollary that every pca has a weak completion (still assuming GCH). The latter statement follows unconditionally from Engeler [6]. GCH is needed in the proof of Theorem 5.1 to guarantee the existence of large cardinals $\kappa$ with $\kappa^{<\kappa}=\kappa$. However, we can get the corollary about completions unconditionally by using a binary version of $\mathcal{K}_{2}^{\kappa}$ defined on $2^{\kappa}$ rather than $\kappa^{\kappa}$. Namely, for the binary version we need the condition $2^{<\kappa}=\kappa$, and ZFC unconditionally proves the existence of arbitrary large $\kappa$ with this property. [^5] Namely, $2^{<\kappa}=\kappa$ holds for any $\kappa$ that is a strong limit, i.e. such that $\lambda<\kappa$ implies $2^{\lambda}<\kappa$, and the strong limit cardinals form a proper class, cf. Jech [11, p50]. This gives an alternative proof of Engeler’s result that every pca is weakly completable.
Note that the condition $\kappa^{<\kappa}=\kappa$ is needed in the definition of $\mathcal{K}_{2}^{\kappa}$ as the analog of the coding of finite sequences $\omega^{<\omega}=\omega$ that is used in the definition of $\mathcal{K}_{2}$. Coding of finite sets also plays a role in the definition of Scott’s graph model $\mathcal{G}$ mentioned above. It is interesting to see how Engeler avoids this condition by explicitly putting the finite sets into his definition of a graph algebra $G(A)$. This has as advantage that no coding is needed. On the negative side, the elements of $G(A)$ are somewhat contrived, and much more complicated than the elements of $\omega^{\omega}$ or $\kappa^{\kappa}$.
Acknowledgments
We thank Henk Barendregt for pointing out that it follows from Engeler [6] that every pca has a weak completion, and for sharing the supplementary note Engeler [7]. We also acknowledge the helpful remarks of a referee that corrected an error in the proof of Lemma 4.3.